[Brahms-l] Shift report 20050621 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Jun 21 2005 - 23:57:03 EDT
Shifter leader was AG
Shifters were 


1. Beam:
   - Physics Beam for ~7 hours
   - Polarization at beginning B/Y: -43.32 +/- 2.28, -46.51 +/- 2.31
   - New polarization (19.20)  B/Y: -44.99 +/- 2.39, -53.25 +/- 2.29
   - New polarization (22.13)  B/Y: -50.15 +/- 2.42, -49.86 +/- 2.28

2. Data: 
   - Runs 15710 - 15723
   - Trig2(1.1M),Trig3(192k),Trig4(176k),Trig5(1.1M)
   - Pedestal run done and installed in tpcped, Non pedestal run done.

3. Problems:
   - The IoServer crashed at 19.52 in run #15716.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jun 21 23:57:28 2005

This archive was generated by hypermail 2.1.8 : Tue Jun 21 2005 - 23:57:41 EDT